# The Importance of Directional Feedback for LLM-based Optimizers

### v1

@@ 1 Introduction
As we show, the presence or absence of directional feedback and the possibility to access them is crucial for LLM-based optimization. For a systematic study of factors that make an optimization process challenging, we choose one of the difficult tasks proposed in the work on OPRO (listed as failure cases in Appendix A of OPRO (Yang et al., 2023)) – navigating a bumpy loss function landscape. We discover that an LLM-based optimizer agent’s performance varies with the information the feedback carries, and, given proper feedback, LLMs can strategically improve over past outputs, which makes this previously unsolvable task solvable. In addition, we demonstrate that using LLMs to “synthesize” feedback from a history of observations and prompts can help optimization too.
@@ 3.1 Fundamentals of LLM Optimization
Similar to Newton’s Method using finite differences as an approximate gradient, perhaps the LLM can implicitly compute a “function”:
@@
If we think this is possible, then the input to the LLM can be tuples of $(p_{\text{tunable}}^{1},r_{1}),...,(p_{\text{tunable}}^{k},r_{k})$. Although it is unclear if this is truly the case, this shows that the optimizer needs to retain a history of how past prompts $p$ have caused the reward $r$.
@@
Unlike directional and non-directional feedback, which usually contains information about how to change $p_{\text{tunable}}$, score-based feedback only gives back a numerical value indicating how well $p_{\text{tunable}}$ performs. In our setup, this means LLM-Optimizer only observes reward $r$ without $f$. This is often referred to as 0th-order feedback.
@@ 3.2 Sequential Prompt Optimization
Inspired by the descent method in numerical optimization, we propose an algorithm that aims to satisfy the requirement of descent methods such that we can reach the extremum. We define the following optimization loop with an LLM-based agent $\pi$. Agent $\pi$ with an initial tunable prompt $p_{\text{tunable}}$ takes a task description $p_{\text{task}}$ from the environment and samples an output $o_{1}$. The environment returns a reward $r_{1}$ and feedback $f_{1}$. An LLM-optimizer stores $(o_{1},r_{1},f_{1},p_{\text{tunable}})$ in a history buffer $\mathbb{H}$. When the buffer becomes large, we subsample $H$ from $\mathbb{H}$. The LLM-optimizer proposes a new tunable parameter $p_{\text{tunable}}^{\prime}$. We make an explicit decision on whether to replace $p_{\text{tunable}}$ with $p_{\text{tunable}}^{\prime}$ based on the reward evaluated on the distribution $o$ and $o^{\prime}$. The full procedure is in Algorithm 1. We now describe the implementation choices for each component of our iterative solver.
Policy
Policy is an LLM that takes in a tunable instruction prompt $p_{\text{tunable}}$, description of task $p_{\text{task}}$ and produces an output $o$. It does not see the history of interaction it did with the environment. This design decision prohibits the policy from controlling its own prompts. It is different from some other existing work. For example, Voyager (Wang et al., 2023) would allow the policy to see all of its interaction histories (and errors they make). React (Yao et al., 2023) also allows the policy to see the full error trace. Allowing the policy to see its past errors is a specific design choice on $p_{\text{tunable}}$. We let the optimizer decide what $p_{\text{tunable}}$ should be (and its decision space includes the human-engineered choices in prior works).

### v2

@@ 1 Introduction
As we show, the presence or absence of directional feedback and the possibility to access them is crucial for LLM-based optimization. For a systematic study of factors that make an optimization process challenging, we choose one of the difficult tasks proposed in the work on OPRO (listed as failure cases in Appendix A of OPRO (Yang et al., 2023)) – navigating a bumpy loss function landscape. We discover that an LLM-based optimizer’s performance varies with the type of information the feedback carries, and, given proper feedback, LLMs can strategically improve over past outputs, which makes this previously unsolvable task solvable. In addition, we demonstrate that using LLMs to “synthesize” feedback from a history of observations and prompts can help optimization too.
@@ 3.1 Fundamentals of LLM Optimization
Similar to Newton’s Method using finite differences as an approximate gradient, perhaps the LLM can implicitly compute the following function:
@@
If we think this is possible, then the input to the LLM can be tuples of $(p_{\text{tunable}}^{1},r_{1}),...,(p_{\text{tunable}}^{k},r_{k})$. Although it is unclear if this is truly the case, this shows that the optimizer needs to retain a history of how past prompts $p$ have influenced the reward $r$.
@@
Unlike directional and non-directional feedback, which usually contains information about how to change $p_{\text{tunable}}$, score-based feedback only gives back a numerical value indicating how well $p_{\text{tunable}}$ performs. In our setup, this means LLM-Optimizer only observes reward $r$ without $f$.
@@ 3.2 Sequential Prompt Optimization
Inspired by the descent method in numerical optimization, we propose an algorithm that aims to satisfy the requirement of descent methods such that we can reach the extremum. We define the following optimization process with an LLM-based agent $\pi$. With an initial tunable prompt $p_{\text{tunable}}$ and a task description $p_{\text{task}}$ from the environment, agent $\pi$ samples an output $o_{1}$. The environment returns a reward $r_{1}$ and feedback $f_{1}$. An LLM-optimizer stores $(o_{1},r_{1},f_{1},p_{\text{tunable}})$ in a history buffer $\mathbb{H}$. When the buffer becomes large, we subsample $H$ from $\mathbb{H}$. The LLM-optimizer proposes a new tunable parameter $p_{\text{tunable}}^{\prime}$. We make an explicit decision on whether to replace $p_{\text{tunable}}$ with $p_{\text{tunable}}^{\prime}$ based on the reward evaluated on the distribution $o$ and $o^{\prime}$. The full procedure is in Algorithm 1. We describe the implementation choices for each component of our iterative solver below.
Agent
Agent is an LLM that takes in a tunable instruction prompt $p_{\text{tunable}}$, description of task $p_{\text{task}}$ and produces an output $o$. Unlike previous work, it does not have access to the history of interactions it had with the environment. This design decision gives the optimizer a higher degree of control over the agent’s behavior. Whether the agent should have access to its own history or what history to access should be determined by the optimizer, not by itself. It is different from some other existing work. For example, Voyager (Wang et al., 2023) would allow the agent to see all of its interaction histories (and errors they make). React (Yao et al., 2023) also allows the agent to see the full error trace. Allowing the agent to see its past errors is a specific design choice on $p_{\text{tunable}}$. We let the optimizer decide what $p_{\text{tunable}}$ should be (and its decision space includes the human-engineered choices in prior works).